Please solve for r
2^(r+5) = 2^(2r-1)

Answer by stanbon(48535)

Please solve for r
2^(r+5) = 2^(2r-1)
---------------
The bases are the same so the exponents must be equal:
Please solve for r
2^(r+5) = 2^(2r-1)
r+5 = 2r-1
r = 6
----------
Cheers,
Stan H.